OpenStudy (anonymous):

What is the solution to this system of linear equations? y = x + 2 y = -3

OpenStudy (lalaly):

is there an x after the -3?

OpenStudy (anonymous):

no there isn't o:

OpenStudy (lalaly):

oh then just substitute the value of y in the above equation to find x

OpenStudy (anonymous):

ah, alright. Thank you!

OpenStudy (lalaly):

:) if u want post ur answer here when u find it

OpenStudy (anonymous):

would the answer be (-1, -3) ?

OpenStudy (anonymous):

yay! :) thanks for the help!

OpenStudy (lalaly):

ur welcome:D:D

OpenStudy (anonymous):

(-5, -3) !!!

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

y = x + 2 y = -3 substituting the vale: -3 = x + 2 x=-3-2=-5

OpenStudy (lgbasallote):

\(\Large \mathsf{\text{he has a point...}}\)

OpenStudy (anonymous):

now I'm confused hah

OpenStudy (anonymous):

just take my answer...:)

OpenStudy (anonymous):

alright lol :o

OpenStudy (lgbasallote):

\(\Large \mathfrak{\text{....or she}}\) what are you @myko ?? btw..i agree with him/her

OpenStudy (anonymous):

i am myko

OpenStudy (anonymous):

:) haha thank you

OpenStudy (lgbasallote):

\(\Large \color{red}{\mathbf{L}}\) \(\large \color{green}{\mathcal{O}}\) \(\huge \color{blue}{\mathbf{L}}\) @myko :P